OpenStudy (vshiroky):

Find the equation of the line that has the given slope and passes through the given point. m=1/2, (2,-2)

OpenStudy (texaschic101):

there are two ways to do this....either use y = mx + b or use y - y1 = m(x - x1). I prefer y = mx + b myself, because to me it is easier. y = mx + b slope(m) = 1/2 (2,-2) x = 2 and y = -2 now we sub...we are looking for b, the y intercept -2 = 1/2(2) + b -2 = 1 + b -2 - 1 = b -3 = b so our equation is : y = 1/2x - 3

OpenStudy (texaschic101):

the other form, point-slope form (y - y1 = m(x - x1)....I only use that form if it specifically asks me to...but that is just me

OpenStudy (vshiroky):

So for like (-3,2) (-6,-7) My answer would be y=3x+11

OpenStudy (texaschic101):

now for that, you fist need to find the slope by using the slope formula : slope = (y2 - y1) / (x2 - x1). And once you find the slope, you use the y = mx + b thing.

OpenStudy (vshiroky):

I think that's what I did

OpenStudy (texaschic101):

oh...I am sorry...I didn't check the answer. If you did that, then your answer could be correct, unless it is asking to put it in standard form

OpenStudy (vshiroky):

It wants slope intercept :)

OpenStudy (the_fizicx99):

Yea its right

OpenStudy (texaschic101):

slope is correct...3 y = mx + b m = 3 (-3,2) x = -3 and y = 2 sub 2 = 3(-3) + b 2 = -9 + b 2 + 9 = b 11 = b yes...your equation is : y = 3x + 11

OpenStudy (vshiroky):

Yay thank you :) I will be posting more questions

OpenStudy (texaschic101):

I wish I could help but I am fixing to leave to go somewhere. I will be back on later.